# Towards flexible perception with visual memory

### v2

@@ 3 Capabilities of a visual memory
Our primary goal is to motivate the concept of a machine visual memory from a variety of different perspectives. To this end, we investigate how such a memory can benefit the following capabilities:
  1. 3.1
    Flexible lifelong learning: adding novel OOD classes
  2. 3.2
    Flexibly trading off compute and memory
  3. 3.3
    Flexibly adding billion-scale data without training
  4. 3.4
    Flexible removal of data: machine unlearning
  5. 3.5
    Flexible data selection: memory pruning
  6. 3.6
    Flexibly increasing dataset granularity
  7. 3.7
    Interpretable & attributable decision-making
@@
3.1 Flexible lifelong learning: adding novel OOD classes (both data and labels)
@@
Standard classifiers, whether trained end-to-end (supervised models) or with a linear classifier (self-supervised models), are not able to handle new information without re-training. For instance, adding new classes or changing labels in an existing model usually involves either re-training or fine-tuning parts of the model. A retrieval-based visual memory, in contrast, is able to process such information in a natural and flexible way, aligning with the requirements of lifelong learning [44]. We tested this by adding data for 64 new classes, along with their new labels, to the visual memory of a pre-trained DinoV2 ViT-L14 model (in addition to the ImageNet train set, which is in-distribution for the model). We took the new classes from the NINCO dataset [45], a dedicated OOD dataset that is designed to have no overlap with existing ImageNet labels and samples. This requires the model to transfer what it has learned to new, unseen concepts. The new task is therefore harder, as the model has to retrieve images from both in-distribution and OOD classes. The resulting visual memory has 1064 classes (1K from ImageNet and 64 from NINCO). Table 1 shows that with a visual memory it is possible to add new classes such that the in-distribution accuracy is maintained without catastrophic forgetting (the new classes only change ImageNet validation performance by 0.02–0.04% depending on the aggregation method), while at the same time reaching very high accuracy on the new OOD classes (approx. 87% top-1) without any training. Figure 11 in the appendix confirms that the samples are indeed OOD for the model, as demonstrated by larger distances to nearest neighbors. This highlights that a visual memory is capable of flexibly adding new information—an important capability since the world is not static.
@@
Table 1: Flexible lifelong learning: adding novel OOD classes. A visual memory of DinoV2 ViT-L14 with ImageNet-train (IN-train) as the memory database is able to handle a simple “insert into memory” operation for 64 out-of-distribution classes (data and labels) from the NINCO dataset [45], leading to high performance on the new classes without affecting top-1 ImageNet validation accuracy.
@@ 3.2 Flexibly trading off compute and memory
Next, we turn our attention to studying the scaling behaviour of visual memory with increasing memory model size. We hypothesize that bigger models will be able to attain similar performance as smaller models with lesser amount of visual memory. This is because, all else being equal, a bigger model should be a better featurizer that requires fewer examples in memory to represent different concepts. We empirically study the scaling behaviour of visual memory based retrieval systems in Figure 3(a) using models of different sizes like DinoV2 ViT models of sizes S/14 (21M params), B/14 (86M params), and L/14 (300M params), as well as CLIP ViT models of sizes B/16 and L/14. We plot the top-1 error rate as a function of number of images in visual memory. The plot demonstrates that for each model, the error rate consistently decreases as we increase the visual memory size. Notably, already with a single exemplar per class in memory, ImageNet validation performance is far beyond chance (41% top-1 error for DinoV2 ViT-L14). It also visualizes the possibility of a flexible trade-off between model size and memory size: e.g. for the different DinoV2 models, the S/14, B/14, and L/14 variant achieve similar performance at 1.28M, $\sim$150K, and $\sim$70K memory capacity respectively. In line with [11], this indicates that a smaller model with large memory can match the performance of a larger model with smaller memory.
(a) ImageNet (million-scale)
(b) JFT (billion-scale)
Figure 3: Memory scaling: flexibly trading off compute and memory. ImageNet top-1 validation error decreases systematically as the memory size is increased. (left) Million-scale memory bank consisting of ImageNet-train labels. (right) Billion-scale memory bank consisting of machine-generated pseudo labels on the JFT dataset [46]. Accuracy continues to decrease even with billion-scale data in memory. The roughly constant offset between models of different sizes suggests the possibility of flexibly trading off compute and memory: The same error rate can be achieved with a small model and large memory, or a large model and a small memory.
@@ 3.3 Flexibly adding billion-scale data without training
Billion-scale dataset with pseudo labels. As demonstrated in Section 3.2, performance systematically improves with increased memory size across both small and large models. We here test how far this trend holds beyond relatively small-scale, well-curated settings like ImageNet-1K by scaling visual memory to the billion-scale unlabeled data regime. We study this by obtaining a large-scale dataset from the union of the ImageNet-1K train set and a subset of the JFT-3B dataset [46]. To this end, we treat JFT as an unlabeled dataset by ignoring its original labels and instead obtaining pseudo labels by running them through ViT-22B-224px [47], a highly performant classifier. We excluded images whose labels do not have a correspondence with the ImageNet labels.
Scaling. In Figure 3(b), we show the downstream performance of two DinoV2-ViT model variants on ImageNet validation examples as a function of visual memory size. The plot demonstrates that even in the billion-scale data regime, validation error decreases when increasing memory size without any training. The gain from more data is most prominent when having fewer samples in memory (e.g., going from 1 to 10 samples per class). In log-log space, a logarithmic function fits the empirical scaling trend well. In the literature, simple scaling trends such as the one we observe are powerful predictors of scaling behaviour for different model and dataset sizes [48; 49; 50].
Out-of-distribution performance. In order to understand whether the benefits of increased memory size transfer to out-of-distribution (OOD) data, we compared DinoV2 ViT-L14 once with ImageNet-train in memory and once with JFT pseudo-labels in memory. The models are evaluated on the ImageNet-A [51], ImageNet-R [52], ImageNet-Sketch [53], ImageNet-V2 [54], and ImageNet-ReaL [55] datasets. As an additional well-performing yet “inflexible” baseline, we report linear probing accuracies from the DinoV2 paper [38]. Table 2 shows that visual memory scaled with JFT data improves OOD performance across all datasets compared to an ImageNet-based visual memory. Gemini re-ranking again improves leads to performance gains. Overall, the finding that memory scale transfers to OOD improvements is important in the context of continual learning, where a flexible visual memory can easily incorporate newly available data that the model was not trained on and improve performance both in- and out-of-distribution.
Table 2: OOD evaluation. Out-of-distribution performance improves with larger visual memory size. Across all OOD datastes, a visual memory of DinoV2 ViT-L14 with JFT-3B (JFT memory) outperforms ImageNet memory demonstrating advantages of scaling visual memory for OOD performance.
@@ 3.4 Flexible removal of data: machine unlearning
The world is not static. Thus, in addition to the need to flexibly add novel data, there is often a need to remove the influence of specific training data from a model’s decision-making process after it has been trained [3; 4; 5; 6]. A range of intricate methods are being developed to remove or reduce the influence of certain training samples [56; 57; 58; 59; 60]—a challenging endeavour if knowledge is embedded in millions or billions of model weights. In contrast, for models with an explicit visual memory, machine unlearning becomes as simple as removing the dataset sample from the visual memory. For instance, after adding the NINCO dataset [45] into visual memory, we can remove any NINCO sample with outstanding performance on all three key unlearning metrics reported by [61]: Efficiency: How fast is the algorithm compared to re-training? (Lightning fast.) Model utility: Do we harm performance on the retain data or orthogonal tasks? (Not at all.) Forgetting quality: How much and how well are the ‘forget data’ actually unlearned? (Completely and entirely.) Can machine unlearning therefore be solved with a visual memory? If the embedding model is trained on data that needs to be unlearned, machine unlearning remains challenging. If, however, the embedding model is trained on a safe, generalist dataset (e.g., a publicly available image dataset) and data that may need to be considered for unlearning later is simply put into the visual memory, then machine unlearning indeed becomes as simple as deleting a datapoint from the visual memory. This can be particularly helpful for tasks that may require private or confidential data—a model can be trained on publicly available datasets to learn general and information features and the private data can be added to a visual memory on local devices for downstream tasks to preserve privacy.
@@ 3.5 Flexible data selection: memory pruning
The ability to flexibly remove the influence of certain datapoints is not just desirable in the unlearning sense, but also advantageous in the context of dataset pruning, an emerging field that analyzes the quality of individual data points. The goal of dataset pruning is to retain only useful samples, while removing those that have a neutral or harmful effect on model quality. The key challenge is that in standard black-box models, it is entirely unclear whether any given sample is helpful or harmful. The gold standard is leave-one-out-training (for ImageNet, this would consist of training 1.28 million models); current methods seek to approximate this extremely costly approach with various heuristics [62; 63; 64; 65; 66]. By contrast, the contribution of a data sample to decision-making in a visual memory based system is straightforward. For any given query image $\tilde{\boldsymbol{x}}$, the neighbor set $\mathsf{Neighbors}(\tilde{\boldsymbol{x}})$ clearly reveals which samples contributed to the decision. Furthermore, this information also highlights whether the samples were helpful (correct label) or harmful (wrong label) for the decision. We, therefore, transfer the concept of dataset pruning to memory, and propose visual memory pruning. To this end, we estimate sample quality by querying the ImageNet training set against a visual memory consisting of the exact same dataset (IN-train, discarding the first neighbor which is identical to the query). This approach requires no more compute than a single forward pass over the training set. We then record the number of times any given neighbor contributed to a wrong decision, resulting in a sample quality estimate. This enables us to exclude low-quality neighbors from the decision-making process by either removing them from the visual memory entirely (“hard memory pruning”) or by reducing their weight compared to higher-quality neighbors (“soft memory pruning”). Method details can be found in Appendix G. In Table 3, we show that both memory pruning variants improve ImageNet validation accuracy, with soft pruning leading to larger gains than hard pruning. Figure 4 visualizes the decision-making process for a randomly selected sample where estimating sample reliability improves decision quality. Given that observing the outcome of an intervention is many orders of magnitude faster in visual memory models (as opposed to traditional leave-out-training), we are optimistic that the visual memory pruning gains we observed with two simple strategies can be improved further in the future.
@@
soft pruning (ours) 83.6 83.6 83.9 84.1
3.6 Flexibly increasing dataset granularity
In contrast to static classification, where a model is trained once without further updates, a visual memory model should be able to flexibly refine its visual understanding as more information becomes available. We test this using DinoV2 ViT-L14 embeddings on the iNaturalist21 dataset [67], a large-scale imbalanced dataset of animal and plant images containing 10,000 species spanning seven taxonomic levels, from coarse (kingdom) to fine-grained (species). In a leave-one-out fashion, we simulate the discovery of a new species by putting 50 exemplars for each of the 9,999 species into memory and then step by step adding more data for the remaining “newly discovered” species—starting from zero exemplars all the way to 50 exemplars (see Algorithm 1 for an algorithmic description). In Figure 5 we observe the following: (1.) Already before a single example of the new species is added, it can already be placed in the right part of the taxonomic tree well beyond chance (35.2% accuracy at the genus level compared to $\sim$0chance accuracy). (2.) Accuracy at the species level improves substantially by adding just a handful of images of the target species (e.g., 5–10 images); a regime where training a classifier would typically fail due to data scarcity. (3.) Interestingly, adding more samples of the discovered species not only improves species-level accuracy, but also leads to a “rising tide lift” of improvements across all levels of the taxonomic hierarchy. Overall, this indicates that a visual memory is well-suited for hierarchical classification tasks and settings where data for new concepts is initially scarce but becomes more abundant over time—which is often the case in applications like fraud detection, personalized recommender systems, and scientific discovery.
Figure 5: Impact of memory bank size on top-1 accuracy across taxonomic levels on iNaturalist. Top-1 accuracy for a target species across different taxonomic levels as the number of exemplars in the memory bank for that species increases from 0 to 50. Each line represents the average accuracy over all 10,000 species in the iNaturalist 2021 dataset, while the number of examples in visual memory is fixed at 50 exemplars for all other species. The black dotted line indicates baseline accuracy from predicting the majority class.
@@ 3.7 Interpretable & attributable decision-making
Attributing model decisions to training data.
Unlike a black-box deep learning model, a visual memory system offers a natural way to understand a model’s specific predictions by attributing them to training data samples [e.g. 27]. In Figure 6, we visualize misclassified validation set examples from the ImageNet-A dataset [51] using a memory of the ImageNet-1K training set. These randomly selected samples illustrate that many seemingly strange errors (e.g., predicting a type of fence instead of a teddy bear, or a unicycle instead of a bow tie) do in fact appear sensible given the data, raising questions about label quality of ImageNet-A—in a similar vein as label issues identified for ImageNet [55; 54; 68]—rather than about model quality.
Understanding compositionality of representations.
Figure 6: Interpretable decision-making. A retrieval-based visual memory enables a clear visual understanding of why a model makes a certain prediction. Here, we show four randomly selected misclassified query images from ImageNet-A [51] along with five nearest neighbors from DinoV2 ViT-L14 using the ImageNet-1K training set as visual memory. All labels are from the respective datasets (ImageNet-A for query and ImageNet-train for neighbors). While all neighbors visually look reasonable, not all labels do.

### v3

@@ 3 Capabilities of a visual memory
Our primary goal is to motivate the concept of a machine visual memory from a variety of different perspectives. To this end, we investigate how such a memory can benefit the following capabilities: 3.1 Flexible lifelong learning: adding novel OOD classes; 3.2 Flexibly trading off compute and memory; 3.3 Flexibly adding billion-scale data without training; 3.4 Flexible removal of data: machine unlearning; 3.5 Flexibly increasing dataset granularity; 3.6 Flexible data selection: memory pruning; 3.7 Interpretable & attributable decision-making.
(a) ImageNet (million-scale)
(b) JFT (billion-scale)
Figure 3: Memory scaling: flexibly trading off compute and memory. ImageNet top-1 validation error decreases systematically as the memory size is increased (i.e., recognition accuracy increases with scale). (left) Million-scale memory consisting of ImageNet-train labels. (right) Billion-scale memory bank consisting of machine-generated pseudo labels on the JFT dataset (Zhai et al., 2022). Accuracy continues to decrease even with billion-scale data in memory. The roughly constant offset between models of different sizes suggests the possibility of a flexible trade-off: The same error rate can be achieved with a small model and large memory, or a large model and a small memory.
@@
3.1 Flexible lifelong learning: adding novel OOD classes (data and labels)
@@
Standard classifiers, whether trained end-to-end (supervised models) or with a linear classifier (self-supervised models), are not able to handle new information without re-training. For instance, adding new classes or changing labels in an existing model usually involves either re-training or fine-tuning parts of the model. A retrieval-based visual memory, in contrast, is able to process such information in a natural and flexible way, aligning with the requirements of lifelong learning (Parisi et al., 2019). We tested this by adding data for 64 new classes, along with their new labels, to the visual memory of a pre-trained DinoV2 ViT-L14 model (in addition to the ImageNet train set, which is in-distribution for the model). We took the new classes from the NINCO dataset (Bitterwolf et al., 2023), a dedicated OOD dataset that is designed to have no overlap with existing ImageNet labels and samples. This requires the model to transfer what it has learned to new, unseen concepts. The new task is therefore harder, as the model has to retrieve images from both in-distribution and OOD classes. The resulting visual memory has 1064 classes (1K from ImageNet and 64 from NINCO). Table 1 shows that with a visual memory it is possible to add new classes such that the in-distribution accuracy is maintained without catastrophic forgetting (the new classes only change ImageNet validation performance by 0.02–0.04% depending on the aggregation method), while at the same time reaching very high accuracy on the new OOD classes (approx. 87% top-1) without any training. Figure 12 in the appendix confirms that the samples are indeed OOD for the model, as demonstrated by larger distances to nearest neighbors. This highlights that a visual memory is capable of flexibly adding new information—an important capability since the world is not static. Furthermore, the memory is highly robust towards label corruption for up to 60% random labels, as shown in Appendix D.
@@
Table 1: Flexible lifelong learning: adding OOD classes. A visual memory of DinoV2 ViT-L14 with ImageNet-train (IN-train) as memory database is able to handle a simple “insert into memory” operation for 64 out-of-distribution classes (data and labels) from the NINCO dataset (Bitterwolf et al., 2023), leading to high performance on new classes without harming existing ones.
@@ 3.2 Flexibly trading off compute and memory
Next, we turn our attention to studying the scaling behaviour of visual memory with increasing memory model size. We specifically test whether smaller models larger memory can match the performance of larger models with smaller memory. This is because, all else being equal, a bigger model might require fewer examples in memory to represent different concepts. We empirically study the scaling behaviour of visual memory based retrieval systems in Figure 3(a) using models of different sizes like DinoV2 ViT models of sizes S/14 (21M params), B/14 (86M params), and L/14 (300M params), as well as CLIP ViT models of sizes B/16 and L/14. We plot the top-1 error rate as a function of number of images in visual memory. The plot demonstrates that for each model, the error rate consistently decreases as we increase the visual memory size. Notably, already with a single exemplar per class in memory, ImageNet validation performance is well above chance (41% top-1 error for DinoV2 ViT-L14). It also suggests the possibility of a flexible trade-off between model size and memory size: e.g. for the different DinoV2 models, the S/14, B/14, and L/14 variant achieve similar performance at 1.28M, $\sim$150K, and $\sim$70K memory capacity respectively. This indicates that a smaller model with large memory can match the performance of a larger model with smaller memory.
@@ 3.3 Flexibly adding billion-scale data without training
Billion-scale dataset with pseudo labels. As demonstrated in Section 3.2, performance systematically improves with increased memory size across both small and large models. We here test how far this trend holds beyond relatively small-scale, well-curated settings like ImageNet-1K by scaling visual memory to the billion-scale unlabeled data regime. Billion-scale search is known to be fast (Johnson et al., 2019; Guo et al., 2020; Chen et al., 2021; Khan et al., 2024), but it is an open question how well a billion-scale memory with a modern featurizer like DinoV2 performs for classification. In order to test predictive performance at scale, we obtain a billion-scale dataset from the union of the ImageNet-1K train set and a subset of the JFT-3B dataset (Zhai et al., 2022). To this end, we treat JFT as an unlabeled dataset by ignoring its original labels and instead obtaining pseudo labels via ViT-22B-224px (Dehghani et al., 2023), a highly performant classifier. We excluded images whose labels do not have a correspondence with ImageNet labels.
Scaling. In Figure 3(b), we show the downstream ImageNet validation performance of two DinoV2-ViTs as a function of memory size. The plot demonstrates that even in the billion-scale data regime, validation error decreases when increasing memory size without any training. In log-log space, a logarithmic function fits the empirical scaling trend well. In the literature, simple scaling trends such as the one we observe are powerful predictors of scaling behaviour for different model and dataset sizes (Hestness et al., 2017; Kaplan et al., 2020; Hoffmann et al., 2022). This experiment confirms that a memory-based system is performant across seven orders of magnitude, improving from both better features and more data.
Out-of-distribution performance. In order to understand whether the benefits of increased memory size transfer to out-of-distribution (OOD) data, we compare DinoV2 ViT-L14 once with ImageNet-train in memory and once with JFT pseudo-labels in memory. The models are evaluated on the ImageNet-A (Hendrycks et al., 2021), ImageNet-R (Hendrycks et al., 2020), ImageNet-Sketch (Wang et al., 2019a), ImageNet-V2 (Shankar et al., 2020), and ImageNet-ReaL (Beyer et al., 2020) datasets. As an additional well-performing yet “inflexible” baseline, we report linear probing accuracies from the DinoV2 paper (Oquab et al., 2023). Table 2 shows that visual memory scaled with JFT data improves OOD performance across all datasets compared to an ImageNet-based visual memory. Gemini re-ranking again leads to performance gains. Overall, the finding that memory scale transfers to OOD improvements is important in the context of continual learning, where a flexible visual memory can easily incorporate newly available data that the model was not trained on, improving performance both in- and out-of-distribution.
Table 2: OOD evaluation. Out-of-distribution performance improves with larger visual memory size. Across all datasets, a visual memory with JFT memory outperforms ImageNet memory demonstrating advantages of scaling visual memory for OOD performance. Probe details: Appendix J.
@@ 3.4 Flexible removal of data: machine unlearning
The world is not static. Thus, in addition to the need to flexibly add novel data, it is often desirable to remove the influence of specific training data from a model’s decision-making process after it has been trained (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022; Zhang et al., 2023). A range of intricate methods are being developed to remove or reduce the influence of certain training samples (Gupta et al., 2021; Sekhari et al., 2021; Ullah et al., 2021; Kurmanji et al., 2024; Sepahvand et al., 2024)—a challenging endeavour if knowledge is embedded in millions or billions of model weights. In contrast, for models with an explicit visual memory and in the context of classification, machine unlearning becomes as simple as removing the dataset sample from the visual memory. For instance, after adding the NINCO dataset (Bitterwolf et al., 2023) into visual memory, we can remove any NINCO sample with outstanding performance on all three key unlearning metrics reported by Liu (2024): Efficiency: How fast is the algorithm compared to re-training? (Very fast: less than 20 milliseconds for deleting a feature/sample from disk.) Model utility: Do we harm performance on the retain data or orthogonal tasks? (By design not at all.) Forgetting quality: How much and how well are the ‘forget data’ actually unlearned? (Completely and entirely by design.) In comparison, the winner of the NeurIPS 2023 machine unlearning challenge (Triantafillou et al., 2024) still suffers from an accuracy gap of 4%, low forgetting quality, and being relatively inefficient requiring 8 epochs of training.
Can machine unlearning therefore be solved with a visual memory? If the embedding model is trained on data that needs to be unlearned, machine unlearning remains challenging. If, however, the embedding model is trained on a safe, generalist dataset (e.g., a publicly available image dataset) and data that may need to be considered for unlearning later is put into the visual memory, then machine unlearning for classification indeed becomes as simple as deleting a datapoint from the visual memory. This can be particularly helpful for tasks that may require private or confidential data: a model can be trained on publicly available datasets to learn general and information features and the private data can be added to a visual memory on local devices for downstream tasks to preserve privacy.
3.5 Flexibly increasing dataset granularity on iNaturalist
In contrast to static classification, where a model is trained once without updates, a visual memory model should be able to flexibly refine its visual understanding as more information becomes available. We test this using DinoV2 ViT-L14 embeddings on the iNaturalist21 dataset (iNaturalistTeam, 2021), a large-scale imbalanced dataset of animal and plant images containing 10,000 species spanning seven taxonomic levels, from coarse (kingdom) to fine-grained (species). In a leave-one-out fashion, we simulate the discovery of a new species by putting 50 exemplars for each of the 9,999 species into memory and then iteratively adding more data for the remaining “newly discovered” species—starting from zero exemplars all the way to 50 exemplars. Classification is performed using $k=1$ nearest neighbors (no aggregation).
In Figure 4 we observe the following: (1.) Already before a single example of the new species is added, it can already be placed in the right part of the taxonomic tree well beyond chance (35.2% accuracy at the genus level compared to $\sim$0% chance). (2.) Accuracy at the species level improves substantially by adding just a handful of images of the target species (e.g., 5–10 images); a regime where training a classifier would typically fail due to data scarcity. (3.) Interestingly, adding more samples of the discovered species not only improves species-level accuracy, but also leads to a “rising tide lift” of improvements across all levels of the taxonomic hierarchy. This indicates that a visual memory is well-suited for hierarchical classification tasks and settings where data for new concepts is initially scarce but becomes more abundant over time—which is often the case in applications like fraud detection, personalized recommender systems, and scientific discovery.
Figure 4: Impact of memory bank size on top-1, k=1 accuracy across taxonomic levels on iNaturalist. Top-1 accuracy for a target species across different taxonomic levels as the number of exemplars in the memory bank for that species increases from 0 to 50. Each line represents the average accuracy over all 10,000 species in the iNaturalist 2021 dataset, while the number of examples in visual memory is fixed at 50 exemplars for all other species. The black dotted line indicates baseline accuracy from predicting the majority class.
@@ 3.6 Flexible data selection: memory pruning
[…]
@@
soft pruning (ours) 83.6 (+0.4%) 83.6 (+0.3%) 83.9 (+0.3%) 84.1 (+0.5%)
@@ 3.7 Interpretable & attributable decision-making
Unlike a black-box deep learning model, a visual memory offers a natural way to understand a model’s specific predictions by attributing them to training data samples (e.g. Papernot & McDaniel, 2018). In Figure 6, we visualize misclassified validation set examples from the ImageNet-A dataset (Hendrycks et al., 2021) using a memory of the ImageNet-1K training set. These randomly selected samples illustrate that many seemingly strange errors (e.g., predicting a type of fence instead of a teddy bear, or a unicycle instead of a bow tie) are in fact sensible given the data, raising questions about label quality of ImageNet-A—in a similar vein as label issues identified for ImageNet (Beyer et al., 2020; Shankar et al., 2020; Yun et al., 2021)—rather than about model quality. This issue is quantified through a human experiment in Appendix N, showing that 2 out of 5 model “errors” on this dataset are instead label errors.
In a separate human experiment (described in Appendix L), we tested whether access to five nearest neighbors helps humans predict model decisions. In this experiment, for each image we provided four label choices including the ground truth label and the model predicted label (if different). The remaining labels were plausible alternatives based on top CLIP predictions for the ImageNet-A test image. For a black-box classifier (no access to neighbors), human accuracy was 56%; this accuracy improved to 83% when given access to four nearest neighbours from our visual memory with a DinoV2 ViT-L14 featurizer, providing strong and falsifiable evidence for the improved interpretability of a visual memory system.
Figure 6: Interpretable decision-making. A visual memory enables a clear visual understanding of why a model makes a certain prediction. Here, we show four randomly selected misclassified query images from ImageNet-A (Hendrycks et al., 2021) with five nearest neighbors from DinoV2 ViT-L14 using the ImageNet-1K training set as memory. Labels are from the respective datasets (query: ImageNet-A; neighbors: ImageNet-train). Visually, all neighbors appear reasonable, but not all labels do.